# Accelerating Containerized Service Delivery at the Network Edge

### v1

@@ IV Evaluation
We conduct extensive experimental evaluations in two environments: a Docker Compose-based emulation on a high-performance x86_64 host (Intel Xeon Silver 4214) and physical edge devices (Raspberry Pi 4 Model B). The emulation enables the simulation of large-scale environments, while the physical setup validates the practical feasibility of PeerSync in real-world scenarios. We select a range of popular AL/ML and commonly used container images of varying sizes, from small base images to large language model (LLM) containers, as listed in Table IV. These images represent diverse use cases, allowing us to evaluate system performance under varying workloads and conditions.
We evaluated PeerSync against two popular container image distribution systems: Dragonfly [2] and Kraken [43], as well as the conventional HTTP-based pull method, referred to as the Baseline. Dragonfly, hosted by the Cloud Native Computing Foundation (CNCF) as an Incubating Level Project, [^4] https://d7y.io/ leverages peer-to-peer (P2P) technology to accelerate file distribution and container image delivery. It is designed to enhance efficiency in distributed environments by reducing reliance on centralized registries. Kraken, an open-source project developed by Uber and deployed in production since early 2018, [^5] https://github.com/uber/kraken also employs P2P technology for Docker image management, replication, and distribution. Kraken is particularly tailored for hybrid cloud environments, offering decentralized capabilities to optimize large-scale image deployments. Although Starlight [11] also focuses on improving container image delivery, it relies on fundamentally different mechanisms, such as runtime trace collection and dependence on centralized registries. These differences make direct comparisons with PeerSync less relevant (see Sec. II). Consequently, we did not include Starlight in our comparisons.
@@
IV-A Docker Compose-based Emulation
IV-A1 Experimental Setup
We create 10 Linux bridge networks ($n$}, where $n=1$ to 101010) to emulate a distributed environment with interconnected routers. Each network contained a router for inter-network communication and seven worker nodes requesting container images. The routers were configured with bandwidth limits (50 Mbps to 1 Gbps), latency, and packet loss using tc(8) [6], simulating real-world ISP and transit networks. Intra-network bandwidth was set to 1 Gbps with zero packet loss.
To model real-world conditions, we vary model inference request frequencies based on the following equation:
@@ IV-A1 Experimental Setup
where $\text{s}_{i}$ is the image size, and $A$ and $B$ are tunable parameters. Higher values of $A$ and $B$ increased the frequency of image-pulling requests, simulating high-demand scenarios. Background traffic generated by iPerf [4] created additional congestion, reflecting the shared link usage typical in real-world applications. All centralized components, including registries, Dragonfly databases, and Kraken trackers, were placed in net_worker1 to simulate centralized cloud environments.
IV-A2 Image Distribution Time
TABLE V: Average Speedup under high request frequency compared to the Baseline (smaller is better).
Net. Profile Baseline Dragonfly[2] Kraken[43] PeerSync
@@ IV-A2 Image Distribution Time
Fig. 5 illustrates the average container image distribution times under different network conditions, providing critical insights into the importance of local data prioritization. Since the Baseline method shows significant performance degradation as request frequency increases, a logarithmic scale (base 222) is used on the $y$-axis for clarity.
@@
Stable network conditions (see top of Fig. 5). In environments with sufficient and stable network resources, all P2P-based methods, including PeerSync, exhibit similar performance trends as request frequency increases (see Table V). While distribution latency increases slightly with higher request loads, P2P methods consistently outperform the Baseline HTTP approach. This aligns with the known benefits of P2P models, such as efficient data sharing and reduced reliance on centralized registries, which have been extensively validated in data center contexts. The trend is particularly pronounced for large LLM images: the average distribution time roughly quadruples as request frequency increases, highlighting the critical role of P2P technologies in facilitating the deployment of large modern AI/ML models with billions of parameters.
@@
Congested network conditions (see lower-left of Fig. 5). In environments with bandwidth limitations, packet loss, and high latency, the Baseline method suffers severe performance degradation due to the long-tail effects of centralized pulling. Remote data transfers are particularly affected, often exceeding the default time limit of 1200 seconds, as indicated by the missing points in the green Baseline lines. In contrast, PeerSync and other P2P-based methods maintain better performance by distributing the load across multiple peers, effectively mitigating bandwidth constraints. Notably, Dragonfly shows a sharp performance drop for larger images due to its reliance on centralized components such as schedulers and managers. As communication between edge nodes and these components becomes unreliable, Dragonfly’s ability to coordinate cloud-edge interactions weakens, resulting in bottlenecks during image serving. PeerSync, by contrast, leverages its decentralized architecture to achieve consistent performance, even under congestion.
@@
Variable network conditions (see lower-right of Fig. 5). This scenario simulates real-world edge environments where latency, packet loss, and bandwidth fluctuate significantly, and participating nodes frequently join and leave the network. Such instability disrupts inter-peer connections and hinders image fetching, even for smaller images like langchain/langchain. Compared to other P2P-based solutions, PeerSync demonstrates superior resilience under these conditions. By prioritizing local peer synchronization and minimizing dependence on centralized components, PeerSync reduces reliance on specific, potentially unstable links and dynamically explores all available links to maximize performance. This strategy not only improves distribution efficiency but also enhances fault tolerance, which is critical in environments with unpredictable network behavior.
@@
PeerSync’s ability to optimize local storage and peer communication makes it particularly robust in dynamic edge environments. As shown in Table V, PeerSync outperforms the Baseline, Dragonfly, and Kraken under unstable, high-frequency scenarios by average factors of 2.72, 1.79, and 1.28, respectively. [^6] In our experiments, the time limit for each download request was set to 1200 seconds. In some instances, the Baseline was unable to finish within this time frame, resulting in missing data points when calculating the average speedup. We assigned a value of 1200 seconds to the missing data points. Consequently, the actual speedup relative to the Baseline is higher than the calculated value. The speedup is even more pronounced for larger container images, reinforcing PeerSync’s effectiveness in addressing the challenges of large-scale image distribution.
IV-A3 Uplink Occupancy
@@
Tables VI, VII, and VIII illustrate the cross-network traffic for each image distribution method across three distinct network profiles: stable, congested, and unstable conditions. Cross-network traffic represents data that traverses routers and encounters bandwidth limits, latency, and packet loss, effectively simulating the challenges of data traversing upstream ISP transit networks and the Internet. Unlike local network bandwidth, which is often abundant, cross-network bandwidth is typically constrained by external factors such as physical link limitations and tariff plans. Thus, minimizing cross-network traffic is critical, particularly in edge computing, where the goal is to prioritize intra-LAN resources as much as possible.
TABLE VI: Sum of cross-network traffic under sufficient and stable network conditions.
@@
Sufficient and stable network conditions. As shown in Table VI, PeerSync consistently minimizes cross-network bandwidth usage under stable conditions. By prioritizing local peer data, PeerSync achieves an average bandwidth consumption of only 0.82 Gbps, significantly lower than Kraken (0.93 Gbps), Dragonfly (1.37 Gbps), and the Baseline (6.18 Gbps). While Dragonfly and Kraken leverage P2P techniques, their reliance on centralized components for coordination and scheduling results in higher cross-network traffic compared to PeerSync. The Baseline method incurs the highest bandwidth usage due to its centralized pulling mechanism, which does not take advantage of local data sharing.
TABLE VII: Sum of cross-network traffic under congested but stable network conditions.
@@
Congested but stable network conditions. In congested but stable network conditions (Table VII), PeerSync continues to outperform other methods by maintaining an average cross-network traffic of just 0.91 Gbps. The Baseline approach shows the lowest maximum bandwidth (10.36 Gbps), which is a result of its slower and less efficient data retrieval. However, it saturates network capacity with an average bandwidth usage of 9.81 Gbps, underscoring its inefficiency in managing congested networks. Kraken and Dragonfly perform better than the Baseline but still exhibit performance degradation due to their centralized dependencies. Dragonfly, in particular, suffers from its reliance on schedulers and managers, resulting in an average traffic of 4.69 Gbps, over five times higher than PeerSync.
TABLE VIII: Sum of cross-network traffic under congested and unstable (i.e. changing packet loss, latency, and bandwidth) network conditions.
@@
Congested and unstable network conditions. Table VIII highlights the performance of each method under congested and unstable network conditions, characterized by fluctuating packet loss, latency, and bandwidth. PeerSync demonstrates remarkable robustness, maintaining the lowest average cross-network traffic at 0.76 Gbps. Kraken performs slightly worse, with an average traffic of 0.89 Gbps, but still manages to leverage its decentralized architecture to reduce reliance on external networks. Dragonfly, on the other hand, suffers significant performance degradation due to its dependence on centralized coordination, resulting in an average traffic of 4.01 Gbps. The Baseline method experiences the lowest maximum bandwidth usage (8.87 Gbps) but continues to exhibit high average traffic (6.34 Gbps), reflecting its inability to adapt to varying network conditions. This highlights the inefficiency of centralized pulling in handling unstable and dynamic environments.
@@
TABLE IX: Number of edge devices in a LAN and the average image distribution time (over 100 image retrieval requests) provided by LAN cache.
Number of edge devices within a LAN 1 2 3 4 5 6 7 8 9 10
TABLE X: Number of nodes (1 $\sim$ 5) and comparison of the sum of cache space occupied by Cache Cleaner and the LRU policy.
@@
Figure 6: Average image distribution time under congested and unstable (i.e. changing packet loss, latency, and bandwidth) network conditions (log scale). The $x$-axis is the parameter $A$. (Smaller images)
IV-A4 Cache Strategy
We evaluated the Cache Cleaner algorithm by analyzing its effectiveness in managing requests and accelerating image retrieval within a LAN. Table IX shows the relationship between the number of edge devices in a LAN and the average time to download images. Initially, as the number of nodes increases, image distribution times rise slightly due to simultaneous requests to the upstream registry. However, as the LAN becomes more populated, the local cache network is increasingly leveraged, resulting in significantly reduced average retrieval times.
@@ IV-A4 Cache Strategy
Comparison with the LRU policy. To further validate the Cache Cleaner’s performance, we compared its total cache usage against the Least Recently Used (LRU) policy in Table X. Cache Cleaner consistently achieves better space utilization by coordinating cache management across neighboring peers. This approach avoids redundant caching of identical content across multiple nodes, enabling more efficient storage utilization within the LAN. In contrast, LRU operates independently on each node and lacks collaboration, leading to higher total cache consumption. Cache Cleaner’s ability to prioritize fast intra-LAN retrieval not only optimizes storage but also improves overall system efficiency.
@@
Distribution of smaller images. Although our primary focus is on large AI/ML images, smaller, non-AI base images are also widely used in practical setups. To emulate the operation of Docker Hub, which serves popular images in large volumes daily, we selected the 10 most downloaded images from Docker Hub and increased the request frequency by adjusting the parameter $A$. Fig. 6 presents the results, demonstrating PeerSync’s ability to maintain its performance advantage over other P2P-based solutions, even under high-frequency request conditions for smaller images. These results highlight PeerSync’s capability to handle the significant volumes typical of Docker Hub operations, proving its versatility across diverse workloads.
The experiment results validate the efficiency of PeerSync’s Cache Cleaner algorithm in optimizing local caching, reducing redundant external requests, and improving retrieval times within a LAN. Its collaborative caching strategy outperforms traditional policies like LRU, offering better space utilization and accelerated data distribution. Additionally, PeerSync demonstrates its ability to efficiently handle high-frequency requests for smaller images, further reinforcing its robustness and scalability across varied workloads and network conditions. This makes PeerSync an efficient solution for both edge and cloud-based environments.
IV-B Real-World Experiments
@@
We deployed a physical testbed using Raspberry Pi devices to compare PeerSync’s performance against the Baseline, Dragonfly, and Kraken. As shown in Fig. 7, the setup comprises six Raspberry Pi (RPi) 4 Model B devices connected via two layer 2 switches, each with a 1 Gbps link speed. Each LAN contains three RPis, with the two switches connected to a common router. The router is configured to forward packets between the networks while limiting the inter-LAN bandwidth to 100 Mbps. To simplify deployment, static routes are configured on the router. This setup is similar to the Docker Compose-based emulation described earlier, with two key differences: a) Physical hardware and real-world networking scenarios are used; b) The aarch64 version of PeerSync is deployed instead of the amd64 version to accommodate the ARM-based Raspberry Pi architecture.
@@
Benchmark requests were made to agents running on each RPi to measure image distribution times. Table XI presents the 90th and 100th percentiles of distribution times achieved by each system. The results demonstrate that PeerSync significantly outperforms the other systems in real-world conditions.
  • At the 90th percentile (P90), PeerSync achieves an image distribution time of 190.79 seconds, representing a 39.7% improvement over Kraken, a 60.9% improvement over Dragonfly, and a 75.9% improvement over the Baseline.
@@
TABLE XI: The 90th and 99th percentile of image distribution times.
@@
V Related Works
Numerous solutions have been explored to enhance distribution speeds in container image delivery. In this section, we briefly review key developments in this area, covering both container-specific solutions and methodologies applied to other resource types.
Before the era of edge computing, several pioneering works focused on fast resource delivery in data center networks. VDN [36] leverages hierarchical network topology and shared chunks between virtual machine images to achieve a 30–80$\times$ speedup for large virtual machine images under heavy traffic in data center networks. Similarly, VMTorrent [37] improves P2P methods used in live streaming, employing block prioritization, profile-based execution prefetch, on-demand fetch, and decoupling of virtual machine image presentation from the underlying data stream. Experiments show VMTorrent can achieve a 30x speedup over traditional network storage.
Since the rise of Docker, containers have become central to deploying workloads in data centers, and distributing container images has emerged as a key issue. While similar to virtual machine images, containers have unique properties, such as a layered structure and registry-based hosting. DevOps practices demand more automated and integrated workload deployment solutions. FID [22], an early P2P-based approach for fast container image distribution, adapts BitTorrent to provide an integrated distribution system. In industry, Dragonfly [2] and Kraken [43] are prominent P2P container image distribution systems. These methods outperform traditional approaches by utilizing spare bandwidth between clients.
However, these solutions do not explicitly address edge computing environments. While Dragonfly [2] and Kraken [43] are P2P-based, they rely on multiple centralized components, which can degrade service quality in the event of component failure. Edge environments are less stable than traditional cloud facilities, and deploying these centralized components in the cloud requires manual configuration across different edge locations. Additionally, edge devices typically have weaker network capabilities, so unrestricted use of P2P protocols can lead to uplink congestion. EdgePier [8], based on IPFS [9], provides a fully decentralized container image distribution system tailored to edge computing. Similarly, Gazzetti et al. [17] proposed a decentralized solution with managers that compute optimal network topologies.

### v2

@@ V Evaluation
We conduct extensive experiments in two environments: a Docker Compose-based emulation on a high-performance x86_64 host (Intel Xeon Silver 4214) and physical edge devices (Raspberry Pi 4 Model B). The emulation enables the simulation of large-scale environments, while the physical setup validates the practical feasibility of PeerSync. We select a range of popular AI/ML and commonly used applications of varying sizes, from small base images to large language model (LLM) containers (Table IV). These images represent diverse use cases, allowing us to evaluate system performance under varying workloads.
We evaluated PeerSync against two popular container image distribution systems: Dragonfly [7] and Kraken [16], as well as the plain HTTP-based pull method, referred to as the Baseline.
  • Dragonfly, hosted by the CNCF as an Incubating Level Project, leverages P2P technology to accelerate content distribution. It is designed to enhance efficiency in distributed environments by reducing reliance on centralized registries.
  • Kraken, an open-source project developed by Uber and deployed in production since early 2018, also employs P2P technology for Docker image management, replication, and distribution. Kraken is particularly tailored for hybrid cloud environments, offering decentralized capabilities to optimize large-scale image deployments.
Although Starlight [10] also focuses on improving container image delivery, it relies on fundamentally different mechanisms, such as runtime trace collection and dependence on centralized registries. These differences make direct comparisons with PeerSync less relevant (see Sec. II). Consequently, we did not include Starlight in our comparisons.
@@
V-A Docker Compose-based Emulation
V-A1 Experimental Setup
We deployed a Docker Compose-based emulation environment comprising 10 Linux bridge networks (net_worker{nn}, $n=1$ to $10$) to model a distributed system of interconnected edge sites. Each network hosted one router responsible for inter-network communication and seven worker nodes issuing container image pull requests. Routers were configured using tc(8) [48] to impose realistic WAN characteristics, including bandwidth limits (20 Mbps to 500 Mbps), variable latency, and packet loss, mimicking ISP and transit links. In contrast, intra-network communication within each LAN was provisioned with uncapped bandwidth and zero packet loss to reflect typical local cluster conditions. This multi-LAN environment models the Internet, where LANs are edge sites and routers are ISPs. In each edge site, traffic is routed internally without flowing through external ISPs.
To emulate heterogeneous user demand, we modeled image request arrival times using an inverse Poisson process:
@@ V-A1 Experimental Setup
where $\text{s}_{i}$ denotes the size of image $i$, and $A$, $B$ are tunable parameters controlling request intensity. Larger values of $A$ and $B$ correspond to higher request frequencies, simulating peak-load scenarios. Background traffic generated via iPerf [49] introduced additional congestion on shared links, better reflecting real-world multi-tenant network usage. All centralized services, including container registries, Dragonfly metadata databases, and Kraken trackers, were co-located in net_worker1 to emulate a cloud-hosted control plane.
For reproducibility of our evaluation, Table V gives full parameters used in our emulation. These parameters include values for network bandwidth and packet loss, node CPU and RAM allocation, and how we configure PeerSync to balance between network and popularity-aware scoring. For the Varying Network experiment group, the bandwidth and packet loss values are randomly modified within the range specified.
Table V: Key parameters used in the emulation.
Parameter Value
Number of LANs 10
Number of Nodes per LAN 7
Emulation Timespan 1800 seconds
Bandwidth (Internal) Uncapped
Packet Loss (Internal) 0%
Bandwidth (Stable Net.) Uncapped
Packet Loss (Stable Net.) 0%
Bandwidth (Congested Net.) 100 to 500Mbps
Packet Loss (Congested Net.) 0 to 10%
Bandwidth (Varying Net., Min) 20 to 150Mbps
Packet Loss (Varying Net., Min) 0 to 10%
Bandwidth (Varying Net., Max) 100 to 500Mbps
Packet Loss (Varying Net., Max) 30 to 50%
Peer Churn Ratio (Varying Net.) 10%
Environmental Variation Cycle (Varying Net.) 60 seconds
Node CPU Allocation 2 to 4 shared cores
Node RAM Allocation 2 to 4 GiB
PeerSync Scoring Weight $\alpha$ 0.5
PeerSync Scoring Weight $\beta$ 0.5
V-A2 Image Distribution Time
Figure 6 presents the average container image distribution times under different network conditions as a function of request frequency (controlled by parameter $A$). Due to the wide dynamic range, especially under high load, a base-2 logarithmic scale is used on the $y$-axis for clarity.
Table VI: Normalized average completion time (Baseline = 100%, lower is better) under different network profiles.
Profile Baseline Dragonfly [7] Kraken [16] PeerSync
@@ V-A2 Image Distribution Time
Stable network conditions (Fig. 6 (a))
@@
In environments with sufficient and stable network resources, all P2P-based methods, including PeerSync, exhibit similar performance trends as request frequency increases (see Table VI). While distribution latency increases slightly with higher request loads, P2P methods consistently outperform the Baseline HTTP approach. This aligns with the known benefits of P2P models, such as efficient data sharing and reduced reliance on centralized registries that have been extensively validated in data center contexts. The trend is particularly pronounced for large LLM images: the average distribution time roughly quadruples as request frequency increases, highlighting the critical role of P2P technologies in facilitating the deployment of modern AI/ML models with billions of parameters.
Congested network conditions (Fig. 6 (b))
@@
In environments with bandwidth limitations, packet loss, and high latency, the Baseline method suffers severe performance degradation due to the long-tail effects of centralized pulling. Remote data transfers are particularly affected, often exceeding the default time limit of 1200 seconds, as indicated by the missing points in the green Baseline lines. In contrast, PeerSync and other P2P-based methods maintain better performance by distributing the load across multiple peers, effectively mitigating bandwidth constraints. Notably, Dragonfly shows a sharp performance drop for larger images due to its reliance on centralized components such as schedulers and managers. As communication between edge nodes and these components becomes unreliable, Dragonfly’s ability to coordinate cloud-edge interactions weakens, resulting in bottlenecks during image serving. PeerSync, by contrast, leverages its decentralized architecture to achieve consistent performance, even under congestion.
Variable network conditions (Fig. 6 (c))
@@
This scenario simulates real-world edge environments where latency, packet loss, and bandwidth fluctuate significantly, and participating nodes frequently join and leave the network. The corresponding parameters are given in Table V. In each variation cycle, every emulated router adopts a new random configuration within the predefined range, 10% of the currently active peers leave the swarm, and a subset of previously disconnected peers rejoin the swarm. Such instability disrupts inter-peer connections and hinders image fetching, even for smaller images like langchain/langchain. Compared to other P2P-based solutions, PeerSync demonstrates superior resilience under these conditions. By prioritizing local peer synchronization and minimizing dependence on centralized components, PeerSync reduces reliance on specific, potentially unstable links and dynamically explores all available links to maximize performance. This strategy not only improves distribution efficiency but also enhances fault tolerance, which is critical in environments with unpredictable network behavior.
@@
PeerSync’s ability to optimize local storage and peer communication makes it particularly robust in dynamic edge environments. As shown in Table VI, PeerSync outperforms the Baseline, Dragonfly, and Kraken under unstable, high-frequency scenarios by average factors of 2.72, 1.79, and 1.28, respectively. [^1] The timeout for each request was 1200 seconds. In some instances, the Baseline could not finish within the limit. We assigned 1200 seconds to the missing data points to avoid missing points. Consequently, the actual speedup relative to the Baseline is higher than the calculated value. The speedup is even more pronounced for larger container images, reinforcing PeerSync’s effectiveness in addressing the challenges of large-scale image distribution.
V-A3 Uplink Occupancy
@@
Tables VII, IX, and X illustrate the cross-network traffic for each image distribution method across three distinct network profiles: stable, congested, and unstable conditions. Cross-network traffic represents data that traverses routers and encounters bandwidth limits, latency, and packet loss, effectively simulating the challenges of data traversing upstream ISP transit networks and the Internet. Unlike local network bandwidth, which is often abundant, cross-network bandwidth is typically constrained by external factors such as physical link limitations and tariff plans. Thus, minimizing cross-network traffic is critical, particularly in edge computing, where the goal is to prioritize intra-LAN resources as much as possible.
Table VII: Aggregate inter-LAN traffic under Stable profile.
@@
Table VIII: Number of devices in a LAN and the average distribution time (over 100 retrieval requests) served by LAN cache.
Number of edge devices within a LAN 1 2 3 4 5 6 7 8 9 10
  • Sufficient and stable network conditions. As shown in Table VII, PeerSync consistently minimizes cross-network bandwidth usage under stable conditions. By prioritizing local peer data, PeerSync achieves an average bandwidth consumption of only 0.82 Gbps, significantly lower than Kraken (0.93 Gbps), Dragonfly (1.37 Gbps), and the Baseline (6.18 Gbps). While Dragonfly and Kraken leverage P2P techniques, their reliance on centralized components for coordination and scheduling results in higher cross-network traffic. The Baseline method incurs the highest bandwidth usage due to its centralized pulling mechanism, which does not take advantage of local data sharing.
  • Congested but stable network conditions. In congested but stable network conditions (Table IX), PeerSync continues to outperform other methods by maintaining an average cross-network traffic of just 0.91 Gbps. The Baseline approach shows the lowest maximum bandwidth (10.36 Gbps), which is a result of its slower and less efficient data retrieval. However, it saturates network capacity with an average bandwidth usage of 9.81 Gbps, underscoring its inefficiency in managing congested networks. Kraken and Dragonfly perform better than the Baseline but still exhibit performance degradation due to their centralized dependencies. Dragonfly, in particular, suffers from its reliance on schedulers and managers, resulting in an average traffic of 4.69 Gbps, over five times higher than PeerSync.
  • Congested and unstable network conditions. Table X highlights the performance of each method under congested and unstable network conditions, characterized by fluctuating packet loss, latency, and bandwidth. PeerSync demonstrates remarkable robustness, maintaining the lowest average cross-network traffic at 0.76 Gbps. Kraken performs slightly worse, with an average traffic of 0.89 Gbps, but still manages to leverage its decentralized architecture to reduce reliance on external networks. Dragonfly, on the other hand, suffers significant performance degradation due to its dependence on centralized coordination, resulting in an average traffic of 4.01 Gbps. The Baseline method experiences the lowest maximum bandwidth usage (8.87 Gbps) but continues to exhibit high average traffic (6.34 Gbps), reflecting its inability to adapt to varying network conditions. This highlights the inefficiency of centralized pulling in handling unstable and dynamic environments.
Table IX: Aggregate inter-LAN traffic under Congested profile.
@@
Table X: Aggregate inter-LAN traffic under Varying profile.
@@
Table XI: Number of nodes (1 $\sim$ 10) and comparison of the sum of cache space occupied by Cache Cleaner and the LRU policy.
@@
Figure 7: Average image distribution time under Varying profile (log scale). The $x$-axis is the parameter $A$.
V-A4 Cache Strategy
We evaluated the Cache Cleaner algorithm by analyzing its effectiveness in managing requests and accelerating image retrieval within a LAN. Table VIII shows the relationship between the number of edge devices in a LAN and the average time to download images. Initially, as the number of nodes increases, image distribution times rise slightly due to simultaneous requests to the upstream registry. However, as the LAN becomes more populated, the local cache network is increasingly leveraged, resulting in significantly reduced average retrieval times.
@@ V-A4 Cache Strategy
To further validate the Cache Cleaner’s performance, we compared its total cache usage against the Least Recently Used (LRU) policy in Table XI. Cache Cleaner consistently achieves better space utilization in multi-node settings by coordinating cache management across neighboring peers. This approach avoids redundant caching of identical content across multiple nodes, enabling more efficient storage utilization within the LAN. In contrast, LRU operates independently and lacks collaboration, leading to higher total cache consumption. Cache Cleaner’s ability to prioritize fast intra-LAN retrieval not only optimizes storage but also improves overall efficiency.
V-A5 Distribution of smaller images
@@
Although our primary focus is on large AI/ML images, smaller, non-AI base images are also widely used in practical setups. To emulate the operation of Docker Hub, which serves popular images in large volumes daily, we selected the 10 most downloaded images from Docker Hub and increased the request frequency by adjusting the parameter $A$. Fig. 7 presents the results, demonstrating PeerSync’s ability to maintain its performance advantage over other P2P-based solutions, even under high-frequency request conditions for smaller images. These results highlight PeerSync’s capability to handle the significant volumes typical of Docker Hub operations, proving its versatility across diverse workloads.
V-A6 Bursty requests
In addition to the experiment settings that assume a Poisson arrival process, we evaluate PeerSync under a bursty request-arrival scenario, where the request arrival rate temporarily increases. Additionally, we mimic diurnal request patterns by applying bursts of different magnitudes. More specifically, this setting is based on the Congested Network AI/ML image distribution experiment, except that the Poisson rate parameter is varied periodically. The arrival rate increases to $1.5\times$ for 20% of the emulation period, and $3.0\times$ for another 10%, thus periodically applying bursty pressure to the swarm. The evaluation results are given in Fig. 6 (d). Empirical results show that PeerSync is able to maintain consistent high service quality under such bursty and diurnal request arrival patterns.
V-B Real-World Experiments
@@
We deployed a physical testbed using Raspberry Pi devices to compare PeerSync’s performance against the Baseline, Dragonfly, and Kraken. As shown in Fig. 8, the setup comprises six Raspberry Pi (RPi) 4 Model B devices connected via two layer 2 switches, each with a 1 Gbps link speed. Each LAN contains three RPis, with the two switches connected to a common router. The router is configured to forward packets between the networks while limiting the inter-LAN bandwidth to 100 Mbps. To simplify deployment, static routes are configured on the router. This setup is similar to the Docker Compose-based emulation described earlier, with two key differences: (i) Physical hardware and real-world networking scenarios are used; (ii) The aarch64 version of PeerSync is deployed instead of the amd64 version to accommodate the ARM-based Raspberry Pi.
@@
Benchmark requests were made to agents running on each RPi to measure image distribution times. Table XII presents the 90th and 99th percentiles of distribution times achieved by each system. The results demonstrate that PeerSync significantly outperforms the other systems in real-world conditions.
  • At the 90th percentile (P90), PeerSync achieves an image distribution time of 190.79 seconds, representing a 39.7% improvement over Kraken, a 60.9% improvement over Dragonfly, and a 75.9% improvement over Baseline.
@@
Table XII: The 90th and 99th percentile of distribution times.
@@
VI Related Works
Numerous solutions have been explored to enhance delivery speeds for container images. In this section, we briefly review key developments in this area, covering both container-specific works and methodologies for other resource types.
[…]
However, these solutions do not explicitly address edge computing environments. While Dragonfly [7] and Kraken [16] are P2P-based, they rely on multiple centralized components, which can degrade service quality in the event of component failure. Edge environments are less stable than traditional cloud facilities, and deploying these centralized components in the cloud requires manual configuration across different edge locations. Additionally, edge devices typically have weaker network capabilities, so unrestricted use of P2P protocols can lead to uplink congestion. EdgePier [9], based on IPFS [25], provides a fully decentralized container image distribution system tailored to edge computing. Similarly, Gazzetti et al. [52] proposed a decentralized solution with managers that compute optimal network topologies. For storage-restricted edge devices, learning-based intelligent caching effectively reduces further container spawn latency [53, 54, 55]. While we also employ this incentive in our work, we focus more on providing a full-lifecycle image management solution for systematic optimization of container image management. Recent advancements in edge computing offloading, exemplified by systems like FlexSlice [19], employ RL/TD3 techniques to generate optimal scheduling decisions in fluctuating scenarios. In contrast, PeerSync is designed with heuristic approaches that do not involve online learning. This decision is motivated by the significant operational requirements of learning-based models, including access to large training datasets, specialized GPU/NPU, and stable feedback loops that are frequently unavailable in typical edge deployments. Consequently, PeerSync’s use of a sliding window and RTT-based metrics provides a computationally efficient solution that achieves near-optimal locality, making it better suited for latency-sensitive, one-shot pull operations.